OpenStudy (anonymous):

Which of the following are solutions to the quadratic equation? 2x2 + 6x - 10 = x2 + 6

OpenStudy (anonymous):

\[2x^2 + 6x - 10 = x^2 + 6\] \[\rightarrow 2x^2- x^2+ 6x - 10-6 = 0\] \[\rightarrow x^2+ 6x - 16 = 0\] \[\rightarrow x^2+ 8x -2x- 16 = 0\] \[\rightarrow x(x+ 8)-2(x+ 8) = 0 \rightarrow (x+ 8)(x-2) = 0\] either x+8=0 or x-2 =0 i.e. x=-8 or x=2 Hence {-8, 2} is the solution set of the given quadratic equation..

OpenStudy (anonymous):

@youngstar10